# Few-Shot Recognition via Stage-wise Retrieval-Augmented Finetuning

### v2

@@ 4.2 Benchmarking Results and Ablation Studies
SWAT significantly outperforms previous methods. Table 2 shows that simply finetuning on few-shot data readily outperforms the previous state-of-the-art methods without bells and whistles. Moreover, our SWAT surpasses previous methods by over 6% in accuracy. In challenging datasets like Semi-Aves, EuroSAT, Aircraft, our improvements are even more significant (10-25%, cf. Appendix Table 7). These results highlight the efficacy of our SWAT for FSR.
Stage-wise finetuning improves rare class accuracy. Table 3 shows overall improvements, with particularly significant gains on rare classes, are achieved through stage-wise finetuning and data augmentation. These results confirm the effectiveness of classifier retraining in handling imbalanced learning on retrieved data.
@@
Further analyses. We conduct more analyses and provide them in Appendix. We summarize salient results here: properly filtering retrieved data improves performance further on some datasets (Table 9); CutMix [74] slightly outperforms other data augmentation methods with minimal computation overhead (Table 11); the string-match retrieval yields the best overall results, though different retrieval methods may excel on specific datasets (Table 10); initializing the classifier with text embedding before finetuning improves performance over random initialization (Table 12), partly explaining why finetuning on few-shot data does not suffer from overfitting; learning on more retrieved images shows diminishing performance gains (Table 15).
@@ 5 Discussions
Limitations and future work. We note several limitations and future directions. First, although our method achieves state-of-the-art performance by simply using hyperparameters reported in the literature, we believe that they are not optimal for FSR. Future work can explore human-in-the-loop intervention for hyperparameter tuning, e.g., manually constructing a validation set using retrieved data to tune hyperparameters. Second, as our method exploits concept names to retrieve pretrained data, future work can employ language models to generate richer descriptions or use the definitions provided in annotation guidelines for better retrieval. Third, despite our work shows that CutMix is effective in mitigating domain gaps and imbalanced distribution, it is worth exploring novel data augmentation methods to further address these issues. Lastly, although our method jointly solves the imbalance issue and domain gap of the retrieved data, future work can design experiments to decouple these two factors for further analysis and develop novel solutions.
@@ 6 Conclusions
We explore few-shot recognition (FSR) motivated from a data annotation perspective. This encourages one to solve FSR towards higher accuracy with a rigorous setup, unlike contemporary FSR works that focus on parameter-efficient finetuning a pretrained Vision-Language Model (VLM) in a flawed setup, which uses an unrealistically large validation set for hyperparameter tuning. We first examine an embarrassingly simple method that finetunes a VLM’s visual encoder on few-shot data. Surprisingly, it readily outperforms previous FSR methods. Then, we adopt the retrieval-augmented learning strategy, which retrieves VLM’s pretraining examples relevant to the concepts of interest to facilitate learning. We identify two interesting issues in the retrieved data: its imbalanced distributions, and its domain gaps with the few-shot examples. We propose a simple yet effective method, Stage-Wise retrieval-Augmented fineTuning (SWAT), to address the two issues. Across nine benchmark datasets, our SWAT outperforms previous methods by 6% accuracy on average, achieving the state-of-the-art FSR performance.
@@
  • Section A. We provide a detailed summary of datasets used in our experiments.
  • Section B. We provide hyperparameters used in our work.
  • Section C. We report detailed results for each benchmark dataset.
  • Section D. We compare different retrieval and filtering methods.
  • Section E. We compare different data augmentation methods.
  • Section F. We compare different classifier initialization methods and provide detailed results of stage-wise training on each dataset.
  • Section G. We provide details of retrieved images for each dataset and analyze the impact of retrieval size.
  • Section H. We show retraining the classifier does not overfit for more datasets.
  • Section I. We provide our code and instructions.
@@ A Summary of Datasets
We summarize the nine fine-grained datasets used in our experiments in Table 6. Following [35, 57], we sample few-shot data from the official training set and evaluate model performance on the official test set, except for ImageNet where we report performance on its validation set. We repeat each experiment for three times with three random seeds. Note that we strictly follow the validation-free protocol [57] that we do not use any validation data for hyperparameter tuning or model selection.
Table 6: Statistics of nine fine-grained datasets repurposed in our work. We list the number of images in the official training, validation, and test sets for each dataset. The protocol of few-shot recognition samples few-shot data from the official training set; we use them as our train set. We repeat the sampling and training for three times for each method with three random seeds. To evaluate methods, we repurpose their official test set as our test set (except on ImageNet where we use its official validation set as our test set). We benchmark methods on our test sets. Note again that we do not use any validation examples for model selection or hyperparameter tuning; instead, we strictly adhere to the realistic validation-free protocol for few-shot research [57].
@@
B Hyperparameter Setting
We follow previous work in other lines [42, 35, 32] to set hyperparameters in our work. Specifically, for stage-1 end-to-end finetuning of SWAT, we follow suggestions from [32, 67] to use a smaller learning rate (1e-6) for updating the visual encoder and a larger learning rate (1e-4) for the linear classifier. We initialize the classifier weights using the text embedding following [42] (Table 12). For other hyperparameters, we adopt the values reported in [42, 35] which include the AdamW optimizer, a batch size of 32, weight decay of 1e-2, and a cosine-annealing learning rate schedule with 50 warm-up iterations. We do not do early stopping as we strictly follow the validation-free protocol [57]. Instead, we train for 50 epochs. The only exception is for ImageNet, we train for 10 epochs due to the large amount of retrieved data for its 1,000 classes. The temperature factor is learned during the finetuning process with an initial learning rate of 1e-4 and the same cosine-annealing learning rate schedule. For data augmentation, we mix retrieved data with few-shot data using CutMix [74], following [44] to sample the mixing ratio from a uniform distribution ($\alpha=1.0$ for beta distribution) and apply CutMix with a probability of 0.5. Our few-shot finetuning (FSFT) adopts the same set of training recipe.
For stage-2 classifier retraining, we use the same set of hyperparameters and follow the practice in [42] to train for 10 epochs with a fixed temperature 0.01. We initialize the classifier in stage 2 using the learned classifier weights from stage-1 end-to-end finetuning, following [32].
For other methods, we reimplement CrossModal Linear Probing [35] using the same hyperparameters as in stage-2 classifier retraining and train for 50 epochs. We obtain the results of CLAP [57] using OpenCLIP models with its default hyperparameters. For other baseline methods that originally uses an unrealistically large validation set for hyperparameter tuning, we copy their results from [57].
@@ B Hyperparameter Setting
Figure 5: Comparison of SWAT with state-of-the-art zero-shot and few-shot methods. We show that simply finetuning the whole visual encoder on few-shot data (our few-shot finetuning, green line) outperforms previous FSR methods, while finetuning on retrieved data (orange line) underperforms zero-shot methods (e.g., ImageNet, EuroSAT, Food, DTD, and Stanford Cars) due to the large domain gap and imbalanced distributions of retrieved data. Our SWAT (red line) outperforms previous methods by $>$6% w.r.t accuracy over nine datasets, with significant improvements (20-30%) on challenging datasets like Semi-Aves and Aircraft. The results validates the effectiveness of our SWAT in mitigating the domain gap and imbalanced distribution issues. We provide detailed performance on each dataset in Table 7. For Flowers, EuroSAT, DTD, and Stanford Cars datasets, we show that SWAT can be further improved by 1-6% of accuracy with proper filtering on the retrieved data (red dashed line, Table 9).
Table 7: Detailed comparison of our SWAT and few-shot finetuning (FSFT) with state-of-the-art zero-shot and few-shot recognition methods using OpenCLIP ViT/B-32 model. SWAT significantly outperforms previous few-shot recognition methods by 6% across nine datasets. We highlight the best number in bold and underline the second best. Superscripts mark improvements compared to previous state-of-the-art FSR method CLAP [57].
@@
Table 8: Detailed comparison of SWAT with state-of-the-art zero-shot and few-shot recognition methods using OpenCLIP ViT-B/16 model. Results show that SWAT achieves larger performance gains ($\sim$8%) than CLAP [57] with a larger backbone. We highlight the best number in bold and underline the second best. Superscripts mark improvements compared to previous state-of-the-art CLAP [57].
@@ C Detailed Benchmarking Results
We compare our SWAT and few-shot finetuning (FSFT) with prior state-of-the-art zero-shot [20, 42] and few-shot recognition methods [35, 57] using the OpenCLIP ViT-B/32 model in Fig. 5 and list the detailed performance in Table 7. We also include the performance our few-shot finetuning without CutMix. Results show that SWAT outperforms previous FSR methods by $>$6% accuracy over nine datasets, with substantial gains (20-30%) on challenging datasets where prior FSR accuracy [35, 57] was below 50% (e.g., Semi-Aves and Aircraft). Additionally, SWAT with OpenCLIP ViT-B/16 model (Table 8) yields even higher gains of 8% over [57] across nine datasets.
Further analysis. Our experiments show that SWAT underperforms prior state-of-the-art FSR method [57] on DTD and Stanford Cars. We conjecture that this is due to the significant domain gaps in the retrieved data, finetuning on which could hurt model’s performance. This motivates us to apply a filtering technique to remove excessively out-of-domain retrieved data. Indeed, as shown in Table 9, applying proper filtering on the retrieved data significantly boosts the performance of SWAT, allowing it to outperform CLAP [57]. We also find filtering improves SWAT on other datasets, including Flowers and EuroSAT. We mark the improved SWAT’s performance on these datasets in Fig. 5.
Moreover, the improved SWAT still underperforms few-shot finetuning (FSFT) on DTD datasets. We hypothesize that the discrepancy is because of DTD’s strict data collection rules, which includes only images that are almost entirely filled with a texture [11]. In contrast, the retrieved images often have only part of region depicting the texture (Fig. 8 and 9). This suggests future work to explore better retrieval or filtering methods to find images that are better aligned with downstream distribution, e.g., by referring to the data collection rules provided in the data annotation guidelines. We explore different retrieval methods in Section D.
Table 9: Comparison of SWAT’s performance with prior state-of-the-art FSR method CLAP [57]. SWAT underperforms CLAP on DTD and Cars datasets due to the significant domain gaps. However, with proper filtering on retrieved images (by keeping the top-10 retrieved images for each class that are ranked by prompt-to-caption or T2T similarity and discarding others), SWAT outperforms CLAP. We show results of different retrieval size in Table 15. Subscripts mark the performance difference compared with CLAP [57].
@@
Table 10: Comparison of SWAT using different retrieval methods. We conduct experiments on six datasets by first conducting string matching following [42] to download images whose captions contain any of the concepts’ synonyms, then ranking the images using different text (few-shot concepts or database captions) and image (database images or few-shot images) features for selecting the images most relevant to downstream concepts. The top-ranking 500 images for each class are selected for running SWAT with 16 few-shot data. Results show that despite all methods outperforms random sampling by $<$1% in average accuracy, no single method is the best for all datasets. We highlight the best number in bold and underline the second best. We further explore adding a text-to-image filtering before text-to-text ranking to remove noisy images with image-to-FS-concept similarity less than 0.25. Results show that T2I filtering improves SWAT’s performance significantly, especially for DTD dataset (6% improvement). We show examples of T2I filtered images in Fig. 10.
@@ D Analysis of Retrieval and Filtering Methods
Retrieval-augmented learning has been extensively studied for zero-shot recognition [36, 65, 42]. Previous work [36] utilizes text-to-text (T2T) or text-to-image (T2I) similarity to retrieve images relevant to each downstream concept. However, as noted by [42], such similarity-based retrieval requires significant storage for downloading all the source images (e.g., $>$10TB for LAION-400M) and high compute costs for computing image and text features (>250 T4 GPU hours). In addition, [65] points out the challenge of threshold selection in similarity-based retrieval: setting it too high includes irrelevant images, which can negatively impact training. Moreover, the proper threshold varies for different concepts, making it infeasible to search at scale. Given the above limitations, in this study, we adopt the string-matching-based retrieval by [42] in the following steps.
Step 1: we use string matching to retrieve images whose captions contain any of downstream concepts’ synonyms. This circumvents the large storage cost, as now we only need to download the text (60GB for LAION texts) for string matching and then the images with matching captions (50GB for all nine datasets). Additionally, [42] shows that using concept synonyms help retrieve diverse images which benefits the retrieval-augmented learning.
Step 2: we then rank the retrieved images based on prompt-to-caption (T2T) similarities and select the top-ranking 500 images for each downstream concepts. We compare different ranking methods using text and image features in Table 10. The results show that, despite all ranking methods outperform the random sampling, no single ranking method is the best across all datasets. This suggests future work to design retrieval methods customized to each downstream task.
In addition, we follow the practice in the curation of LAION dataset to apply text-to-image (T2I) filtering, excluding noisy retrieved images with T2I (few-shot concepts and retrieved images) similarities below 0.25. Results in Table 10 show that this filtering notably improves SWAT’s performance, especially on DTD dataset ($>$6%). We show examples of filtered noisy images in Fig. 10.

### v3

@@ 4.2 Benchmarking Results and Ablation Studies
SWAT significantly outperforms previous methods. Table 2 shows that simply finetuning on few-shot data readily outperforms the previous state-of-the-art methods without bells and whistles. Moreover, our SWAT surpasses previous methods by over 6% in accuracy. In challenging datasets like Semi-Aves, EuroSAT, and Aircraft, our improvements are even more significant (10-25%, cf. Appendix Table 7). These results highlight the efficacy of our SWAT for FSR.
Stage-wise finetuning improves rare class accuracy. Table 3 shows overall improvements, with particularly significant gains on rare classes, achieved through stage-wise finetuning and data augmentation. These results confirm the effectiveness of classifier retraining in handling imbalanced learning on retrieved data.
@@
Further analyses. We conduct more analyses and provide them in the Appendix. We summarize the salient results here: properly filtering retrieved data improves performance further on some datasets (Table 9); the string-matching-based retrieval yields the best overall results, though different retrieval methods may excel on specific datasets (Table 10); CutMix [76] slightly outperforms other data augmentation methods with minimal computation overhead (Table 11); SWAT outperforms other ensembling-based[69] or contrastive finetuning methods[17] (Table 13); initializing the classifier with text embedding before finetuning surpasses random initialization (Table 14), partly explaining why finetuning on few-shot data does not overfit; learning on more retrieved images shows diminishing performance gains (Table 17).
@@ 5 Discussions
Limitations and Future Work. We note several limitations and future directions. First, although our method achieves state-of-the-art performance by simply using hyperparameters reported in the literature, we believe that they are not necessarily optimal for each FSR task. Future work can explore human-in-the-loop intervention for hyperparameter tuning, e.g., manually constructing a validation set using retrieved data to tune hyperparameters. Second, as our method exploits concept names to retrieve pretrained data, future work can employ language models to generate richer descriptions or use the definitions provided in annotation guidelines for better retrieval. Third, despite our work showing that CutMix is effective in mitigating domain gaps and imbalanced distribution, it is worth exploring novel data augmentation methods to further address these issues. Lastly, although our method jointly solves the imbalance issue and domain gap of the retrieved data, future work can design experiments to decouple these two factors for further analysis and develop novel solutions.
@@ 6 Conclusions
We explore few-shot recognition (FSR) motivated from a data annotation perspective. This encourages one to solve FSR towards higher accuracy with a rigorous setup, unlike contemporary FSR works that focus on parameter-efficient finetuning a pretrained Vision-Language Model (VLM) in a flawed setup, which uses an unrealistically large validation set for hyperparameter tuning. We first examine an embarrassingly simple method that finetunes a VLM’s visual encoder on few-shot data. Surprisingly, it readily outperforms previous methods. Then, we adopt the retrieval-augmented learning strategy, which retrieves VLM’s pretraining examples relevant to the concepts of interest to facilitate learning. We identify two interesting issues in the retrieved data: its imbalanced distributions, and its domain gaps with the few-shot examples. We propose a simple yet effective method, Stage-Wise retrieval-Augmented fineTuning (SWAT), to address these two issues. Across nine benchmark datasets, our SWAT outperforms previous methods by 6% accuracy on average, achieving the state-of-the-art FSR performance.
Acknowledgements
This work was supported by FDCT (0067/2024/ITP2), University of Macau (SRG2023-00044-FST), and the Institute of Collaborative Innovation. Tian Liu acknowledges the support from Prof. James Caverlee and the CSE Department at Texas A&M University. Portions of this research were conducted with the advanced computing resources and consultation provided by Texas A&M High Performance Research Computing.
@@ Acknowledgements
  • Section A. We provide a detailed summary of benchmarking datasets used in our experiments.
  • Section B. We provide details of hyperparameters used in our work.
  • Section C. We report detailed results of comparing SWAT with previous FSR methods for each benchmark dataset.
  • Section D. We provide details on how we retrieve pretraining data and compare different retrieval and filtering methods.
  • Section E. We compare different mixed sample data augmentation methods and analyze the impact of the mixing ratio within a batch.
  • Section F. We validate the design of our SWAT by ablating different stage-2 training strategies and comparing SWAT with recent state-of-the-art finetuning methods.
  • Section G. We provide further analyses on SWAT, including the impact of training epochs, different classifier initialization methods, and more detailed experimental results.
  • Section H. We provide analysis of the imbalance of retrieved data and the impact of retrieval size.
  • Section I. We provide code and instructions for replicating our experiments.
@@ A Summary of Datasets
We summarize the nine fine-grained datasets used in our experiments in Table 6. Following [36, 59], we sample few-shot data from the official training set and evaluate model performance on the official test set, except for ImageNet where we report performance on its validation set. We repeat each experiment three times with three random seeds. Note that we strictly follow the validation-free protocol [59] that we do not use any validation data for hyperparameter tuning or model selection.
Table 6: Statistics of nine fine-grained datasets repurposed in our work. We list the number of images in the official training, validation, and test sets for each dataset. The protocol of few-shot recognition samples few-shot data from the official training set; we use them as our train set. We repeat the sampling and training three times for each method with three random seeds. To evaluate methods, we repurpose their official test set as our test set (except on ImageNet where we use its official validation set as our test set). We benchmark methods on our test sets. Note again that we do not use any validation examples for model selection or hyperparameter tuning; instead, we strictly adhere to the realistic validation-free protocol for few-shot research [59].
@@
Refer to caption Refer to caption Refer to caption
Figure 5: Comparison of SWAT with state-of-the-art zero-shot and few-shot methods. We show that simply finetuning the whole visual encoder on few-shot data (our few-shot finetuning, green line) outperforms previous FSR methods while finetuning on retrieved data (orange line) underperforms zero-shot methods (e.g., ImageNet, EuroSAT, Food, DTD, and Stanford Cars) due to the large domain gap and imbalanced distributions of retrieved data. Our SWAT (red line) outperforms previous methods by $>$6% w.r.t accuracy over nine datasets, with significant improvements (20-30%) on challenging datasets like Semi-Aves and Aircraft. The results validate the effectiveness of our SWAT in mitigating the domain gap and imbalanced distribution issues. We also show that our SWAT+ (red dashed line) which finetunes both visual encoder and classifier on few-shot data in stage 2 improves further over SWAT (cf. Section F). Detailed performance on each dataset is provided in Table 7. For Flowers, EuroSAT, DTD, and Stanford Cars datasets, we show that SWAT can be further improved by 1-6% of accuracy with proper filtering on the retrieved data (cf. Table 9).
B Hyperparameter Setting
Stage-1 End-to-End Finetuning. We follow previous work in other lines [43, 36, 33] to set hyperparameters in our work. Specifically, for stage-1 end-to-end finetuning of SWAT, we follow suggestions from [33, 69] to use a smaller learning rate (1e-6) for updating the visual encoder and a larger learning rate (1e-4) for the linear classifier. We initialize the classifier weights using the text embedding following [43] (cf. Table 14). For other hyperparameters, we adopt the values reported in [43, 36] which include the AdamW optimizer, a batch size of 32, weight decay of 1e-2, and a cosine-annealing learning rate schedule with 50 warm-up iterations. We do not do early stopping as we strictly follow the validation-free protocol [59]. Instead, we train for 50 epochs. The only exception is for ImageNet, we train for 10 epochs due to the large amount of retrieved data for its 1,000 classes. The temperature factor is learned during the finetuning process with an initial learning rate of 1e-4 and the same cosine-annealing learning rate schedule. For data augmentation, we mix retrieved data with few-shot data using CutMix [76], following [45] to sample the mixing ratio from a uniform distribution ($\alpha=1.0$ for beta distribution) and apply CutMix with a probability of 0.5. Our few-shot finetuning (FSFT) adopts the same set of training recipe.
Stage-2 Classifier Retraining. We use the same set of hyperparameters and follow the practice in [43] to train for 10 epochs with a fixed temperature of 0.01. We initialize the classifier in stage 2 using the learned classifier weights from stage-1 end-to-end finetuning, following [33].
Table 7: Detailed comparison of our SWAT and few-shot finetuning (FSFT) with state-of-the-art zero-shot and few-shot recognition methods using OpenCLIP ViT/B-32 model. SWAT significantly outperforms previous few-shot recognition methods by 6% across nine datasets. We also include the results of FSFT with and without CutMix, as well as SWAT+ where we finetuning the whole model rather than only the classifier on few-shot data in the second stage (cf. Section F). We highlight the best number in bold and underline the second best. Superscripts mark improvements compared to previous state-of-the-art FSR method CLAP [59].
@@ B Hyperparameter Setting
SWAT+ (ours) 64.7+5.5 98.3+3.5 60.2+21.1 93.5+11.8 69.8-0.1 92.2+3.8 79.1+0.6 89.2+1.4 69.3+2.2 79.6+6.7
Baselines. For baseline methods, we reimplement CrossModal Linear Probing [36] using the same hyperparameters as in stage-2 classifier retraining and training for 50 epochs. We obtain the results of CLAP [59] using OpenCLIP models with its default hyperparameters. For other baseline methods that originally used an unrealistically large validation set for hyperparameter tuning, we copy their results from [59].
Table 8: Detailed comparison of SWAT with state-of-the-art zero-shot and few-shot recognition methods using OpenCLIP ViT-B/16 model. Results show that SWAT achieves larger performance gains ($\sim$8%) over CLAP [59] with a larger backbone of ViT-B/16. We also include the results of FSFT with and without CutMix, as well as SWAT+ where we finetuning the whole model rather than only the classifier on few-shot data in the second stage (cf. Section F). We highlight the best number in bold and underline the second best. Superscripts mark improvements compared to previous state-of-the-art CLAP [59].
@@ C Detailed Benchmarking Results
We compare our SWAT and few-shot finetuning (FSFT) with prior state-of-the-art zero-shot [21, 43] and few-shot recognition methods [36, 59] using the OpenCLIP ViT-B/32 model in Fig. 5 and list the detailed performance in Table 7. We also include the performance of our few-shot finetuning without CutMix. Results show that SWAT outperforms previous FSR methods by $>$6% accuracy over nine datasets, with substantial gains (20-30%) on challenging datasets where prior FSR accuracy [36, 59] was below 50% (e.g., Semi-Aves and Aircraft). Additionally, SWAT with OpenCLIP ViT-B/16 model (Table 8) yields even higher gains of 8% over [59] across nine datasets.
Further Analysis. Our experiments show that SWAT underperforms prior state-of-the-art FSR method [59] on DTD and Stanford Cars. We conjecture that this is due to the significant domain gaps in the retrieved data, finetuning on which could hurt the model’s performance. This motivates us to apply a filtering technique to remove excessively out-of-domain retrieved data. Indeed, as shown in Table 9, applying proper filtering on the retrieved data significantly boosts the performance of SWAT, allowing it to outperform CLAP [59]. We also find filtering improves SWAT on other datasets, including Semi-Aves, Flowers, and EuroSAT (cf. Table 10).
Moreover, the improved SWAT still underperforms our few-shot finetuning (FSFT) on DTD datasets. We hypothesize that the discrepancy is because of DTD’s strict data collection rules, which include only images that are almost entirely filled with a texture [11]. In contrast, the retrieved images often have only part of the region depicting the texture (Fig. 11 and 12). This suggests future work to explore better retrieval or filtering methods to find images that are better aligned with downstream distribution, e.g., by referring to the data collection rules provided in the data annotation guidelines. We explore different retrieval methods in Section D below.
Table 9: Comparison of SWAT’s performance with prior state-of-the-art FSR method CLAP [59]. SWAT underperforms CLAP on DTD and Cars datasets due to the significant domain gaps. However, with proper filtering on retrieved images (by keeping the top-10 retrieved images for each class that are ranked by prompt-to-caption or T2T similarity and discarding others), SWAT outperforms CLAP. We show results of different retrieval sizes in Table 17. Subscripts mark the performance difference compared with CLAP.
@@
Table 10: Comparison of SWAT using different retrieval methods. We conduct experiments on six datasets by first conducting string matching following [43] to download images whose captions contain any of the concepts’ synonyms, then ranking the images using different text (few-shot concepts or database captions) and image (database images or few-shot images) features for selecting the images most relevant to downstream concepts. The top-ranking 500 images for each class are selected for running SWAT with 16 few-shot data. Results show that despite all methods outperforming random sampling by $<$1% in average accuracy, no single method is the best for all datasets. We highlight the best number in bold and underline the second best. We further explore adding text-to-image filtering before text-to-text ranking to remove noisy images with image-to-FS-concept similarity of less than 0.25. Results show that T2I filtering improves SWAT’s performance significantly, especially for the DTD dataset (6% improvement). We show examples of T2I filtered images in Fig. 13.
@@ D Analysis of Retrieval and Filtering Methods
Retrieval-augmented learning has been extensively studied for zero-shot recognition [37, 67, 43]. Previous work [37] utilizes text-to-text (T2T) or text-to-image (T2I) similarity to retrieve images relevant to each downstream concept. However, as noted by [43], such similarity-based retrieval requires significant storage for downloading all the source images (e.g., $>$10TB for LAION-400M) and high compute costs for computing image and text features (>250 T4 GPU hours). In addition, [67] points out the challenge of threshold selection in similarity-based retrieval: setting it too low includes irrelevant images, which can negatively impact training. Moreover, the proper threshold varies for different concepts, making it infeasible to search at scale. Given the above limitations, in this study, we adopt the string-matching-based retrieval by [43], detailed in the following two steps.
Step 1: String Matching with Synonyms. We use string matching to retrieve images whose captions contain any of the downstream concepts’ synonyms. This circumvents the large storage cost, as now we only need to download the text (60GB for LAION texts) for string matching and then the images with matching captions (50GB for all nine datasets). Additionally, [43] shows that using concept synonyms helps retrieve diverse images which benefits retrieval-augmented learning.
Step 2: Selection by Ranking. To select images that are most relevant to downstream concepts, we rank the retrieved images based on prompt-to-caption (T2T) similarities and select the top-ranking 500 images for each downstream concept. We compare different ranking methods using text (image captions or downstream concepts) and image (pretraining images or few-shot images) features in Table 10. The results show that, despite all ranking methods outperforming the random sampling, no single ranking method is the best across all datasets. This suggests future work to design retrieval methods customized to each downstream task.
T2I Filtering Improves SWAT Performance. To explore better retrieval methods, we follow the practice in the curation of the LAION dataset to apply text-to-image (T2I) filtering, excluding noisy retrieved images with T2I (few-shot concepts and retrieved images) similarities below 0.25. Despite that adding T2I filtering increases the imbalance of retrieved data, it notably improves SWAT’s performance (cf. Table 10), especially on the DTD dataset ($>$6%). We show examples of T2I-filtered noisy images in Fig. 13. This suggests future retrieval methods to explore better filtering techniques. By default, our SWAT does not apply T2I filtering as post-processing, because determining a proper threshold for each class requires a large validation set which is not allowed in our realistic FSR setup.